Tohban Report 2014-10-15
From RHESSI Wiki
|Start Date:||08 Oct 2014|
|End Date:||15 Oct 2014|
|Tohban:||Juan Carlos Martínez Oliveros|
|Next Tohban:||Juan Camilo Buitrago-Casas|
|List all reports|
Solar activity was low this week with some sporadic M class flares. Currently there are 4 active regions on disk and the active region responsible for the latest M class flares will be rotating onto the disk this week.
How many GOES flares occurred?
Flares above B, C, M, X class were 7 18 5 0
And how many of these are listed in the RHESSI flare list?
Flares above B, C, M, X class were 6 9 5 0
And how many had EXCELLENT coverage?
Flares above B, C, M, X class were 5 2 2 0
There were RHESSI flares/GOES flares 127 / 30 over the time range 08-Oct-14 15-Oct-14
The SSR was emptying by the end of the daily pass set during most of the week. A joint observation with SDO made the SSR to reach a maximum of ~44% and as the last BGS contact the SSR fill levels was at about ~9% most likely due to the latest flares.
Cold tip temperatures where a little high this week triggering two warnings. As the last pass the colt tip temperatures are 127.1 K and 125.4 K.
The spacecraft is operating nominally. The attenuators are working correctly. A special operation during SDO rolling maneuver was schedule from 2014/281 17:51 to 282 00:41 during which we intend to collect as much RHESSI/SAS data as possible. Everything worked as planned and additional passes allowed to minimize the impact on the SSR fill level.
The decimation was changed on October 13 at19:59 UT to NORMAL/VIGOROUS and event collection during nights from both the front and rear segments was enable.
No gaps this week.
Detector 2 continue showing light leak issues as well as detector 5.
Detector 9, still show some spikes but the number of spikes became less frequent.
One spike was observed on October 8 in detector 5. http://sprg.ssl.berkeley.edu/~tohban/browser/?show=grth1+qlpcr+qlpfr+qlprr+monfs+monfl+monfr+monff+monrs+monrl+monrr+monrf&date=20141008&time=065450&bar=1
There were no threshold changes this week.
|HLAT Decimation||Rear decimation weight 6, no front decimation|
|Night time data (fronts)||full nighttime events|
|Night time data (rears)||full nighttime events|
|Require extra passes?||Yes|
|Requirement for moving pointer?||No|
|Detector problems?||Detector 5, 9 spikes|